Let z be such that 3(ū + 2) + ū = 2z - ū where ū= (1,3), Ū = (0,4), ū = 4(1, 1). Then ž equals:
Answer #1

Let ² be such that, 3(u + 7) + w = 2 Z - ze 3 u + 3ž to -2 Z -ū Ž = -w-za- Z Now, given that: ū=(1,3), v = (0,4) in = 48(1, 1